OpenStudy (anonymous):

Name the property for the given statement. 2(10 + 15) = 2(10) + 2(15)

rvc (rvc):

i think distributive well im not sure

OpenStudy (anonymous):

its distributive property. distributive deals with opening of brackets a(b+c)=ab+ac 2(10 + 15) = 2(10) + 2(15)

OpenStudy (anonymous):

you there?

OpenStudy (anonymous):

2(10+15)=2*10+2*15

OpenStudy (michele_laino):

I think that property is the distributive property of multiplication over addition

OpenStudy (anonymous):

yeah even i think same but i think there is one thing missing

OpenStudy (anonymous):

nothing is missing if we check the left hand side 2(10+15) =2(25) =50 now for right hand side , we have 2*10+2*15=20+30 =50 then the property is proved because LF=RF